OpenStudy (anonymous):

( Am i right. I think the answer is d.) Suppose E is an event whose probability of occurring is 2/3. What is the probability that E will NOT occur? A) -2/3 B) -3/2 C) 1/2 D) 1/3

OpenStudy (paki):

may be...

jimthompson5910 (jim_thompson5910):

D is correct P(E occurring) + P(E not occurring) = 1 P(E not occurring) = 1 - P(E occurring) P(E not occurring) = 1 - 2/3 P(E not occurring) = 1/3

OpenStudy (anonymous):

awesome just making sure. thanks! :)